height of waterfall is 500m high. of the water retains 65% of the heat generated at the end of the fall , calculate the change in temperature due to the fall

1
Expert's answer
2021-09-01T11:29:36-0400

Water is at height (h)=500m.


putting mass of water= 1kg


Potential energy =mgh="1\times 9.81\times 500=\frac{4905J}{litre}"


"65\% of 4905=\frac{3188.25J}{litre}"


Specific heat of water =4180J"\/" kg"\/" k.


So, E=m"\times" c"\times" ΔT


3188.25="1\times 4180\times" ΔT


ΔT "=0.76^oC"